OpenStudy (anonymous):

Name the line and graph (-1,-2) with a slope of 2/3

OpenStudy (anonymous):

Using y= m*x+b, m is your slope so y= 2/3*x+b, your x and y is here to, so -2 = 2/3*-1+b are you able to work it from here to figure out b?

OpenStudy (anonymous):

Yeah I think so multiply 2/3(-1) to get neg 2/3 then add 2/3 to both sides to get b

OpenStudy (anonymous):

It's actually: -2 -(-2/3) = b because the reverse of -2 = -2/3+b which is?

OpenStudy (anonymous):

-4/3

OpenStudy (anonymous):

correct.

OpenStudy (anonymous):

Now that you know m and b, you can solve any x and y points.